A swimmer heads upstream towards the opposite bank of 200m wide channel in a direction which forms an angle of 60 degree with the bank.at what angle to the bank must the swimmer head in order to reach its destination if the velocity of the boat relative to water is 18km/hr and the flow-velocity of water in the channel is 0.5 m/s ?how long does it take for the swimmer to reach its destination ?for small angles ,it can be assumed that sinA=tanA=Aradian.
1
Expert's answer
2012-06-07T09:05:41-0400
The swimmer's speed V = 18km/hr = 5m/s
The upstream component of swimmer's
speed is Vs = V*cos(60) = 5*0.5 = 2.5m/s.
The other component is Vc =
V*sin(60) = 5*0.8660 = 4.33m/s.
Time needed t = L / Vc = 200 / 4.33 =
46.1894s